# OFER: Occluded Face Expression Reconstruction

### v1

@@ Abstract
Reconstructing 3D face models from a single image is an inherently ill-posed problem, which becomes even more challenging in the presence of occlusions. In addition to fewer available observations, occlusions introduce an extra source of ambiguity, where multiple reconstructions can be equally valid. Despite the ubiquity of the problem, very few methods address its multi-hypothesis nature. In this paper we introduce OFER, a novel approach for single-image 3D face reconstruction that can generate plausible, diverse, and expressive 3D faces, even under strong occlusions. Specifically, we train two diffusion models to generate the shape and expression coefficients of a face parametric model, conditioned on the input image. This approach captures the multi-modal nature of the problem, generating a distribution of solutions as output. Although this addresses the ambiguity problem, the challenge remains to pick the best matching shape to ensure consistency across diverse expressions. To achieve this, we propose a novel ranking mechanism that sorts the outputs of the shape diffusion network based on the predicted shape accuracy scores to select the best match. We evaluate our method using standard benchmarks and introduce CO-545–a new protocol and dataset designed to assess the accuracy of expressive faces under occlusion. Our results show improved performance over occlusion-based methods, with added ability to generate multiple expressions for a given image.
Figure 1: Reconstructions generated with OFER. Our method can reconstruct faces under hard occlusions, providing multiple solutions with diverse expressions that align with the input image. A novel ranking mechanism selects a unique identity from a probabilistic diffusion model, on top of which the expression variants are applied.
@@ 1 Introduction
The task becomes even more challenging when images are subjected to occlusions, as shown in Fig. 1. The main difficulty arises from the face being only partially visible, introducing an additional source of ambiguity– the occluded areas can correspond to an infinite number of valid shapes, making it essentially a multi-hypothesis reconstruction problem. Occlusions are very common in images captured in uncontrolled environments due to factors such as hair, accessories, medical masks, or even strong profile poses and head rotations, as seen in Fig. 9. Despite its prevalence, face reconstruction under such conditions has seldom been addressed.
@@
3D face reconstruction typically involves recovering the parameters of a statistical 3D facial model, such as the 3D Morphable Model (3DMM) [4], either via model fitting [4, 1] or regression [10, 13, 22, 48]. While these works may handle milder occlusions depending on their training data, their performance deteriorates significantly under more severe occlusions (see Fig. 2). A few methods have been specifically proposed to address the challenge of occluded faces [18, 19, 31, 14]. Most offer a unique solution due to their deterministic nature, ignoring the multi-hypothesis nature of the problem. An exception to this is Diverse3D [14], which employs Determinantal Point Process (DPP) [29] to sample a diverse set of outputs. While DDP is designed to capture the diversity of the data, it fails to adequately representthe underlying distribution, often resulting in unrealistic and exaggerated face reconstructions as shown in Fig. 2.
@@
To ensure plausible face reconstructions while addressing the challenges of diverse sampling, we introduce OFER, a novel method for reconstructing 3D faces under occlusions. At its core, OFER employs two denoising diffusion probabilistic models (DDPMs) [26] to generate the shape and expression coefficients of the FLAME [33] parametric face model, conditioned on an input image. The ability of diffusion models to learn the underlying data distribution enables OFER to produce multiple plausible hypotheses. We leverage two pre-trained face image recognition networks as image encoders, and following MICA [55], train our models using only a small dataset of paired 2D-3D data.
@@
Generating diverse solutions is crucial for reconstructing faces under occlusions. However, to produce multiple expressive solutions with a consistent shape structure, it is important to select a unique shape from the set of generated shape coefficients that best represents the face in the image. To achieve this, we propose a novel ranking mechanism that evaluates and ranks those generated samples. Specifically, given an input image along with N generated shape samples, this network scores and ranks these samples, then selects the top-ranked one. To the best of our knowledge, such a ranking mechanism as selection process for diffusion models has not been considered before.
@@
Figure 2: Comparison against occluded Expression Reconstruction methods. We show results from EMOCA [10] (pink); samples generated by Diverse3D [14] (blue); and samples generated by our method (green). EMOCA can only reconstruct a single solution due to its deterministic nature, while Diverse3D shows non-plausible faces. OFER (our method) generates diverse 3D faces with plausible expressions.
@@
To evaluate our approach, we introduce a new dataset and evaluation protocol, CO-545, built from the CoMA dataset [41]. CO-545 includes 545 ground-truth pairs of occluded images and corresponding FLAME non-occluded 3D vertices. Experimental results show that our method effectively generates multiple hypotheses for single in-the-wild occluded images, achieving superior quality and diversity on CO-545. Additionally, when evaluated on NoW benchmark [45], our method demonstrates improved performance with ranking-selection mechanism.
@@ 2 Related Work
3D face reconstruction from a single image has been a research focus for several decades [20], with approaches broadly divided into model-free methods [12, 21, 44, 17, 2, 27, 46, 51] and model-based methods [55, 22, 10, 52, 42]. The latter involves recovering the low-dimensional parameters of a statistical model of the 3D face, e.g., BFM [39] or FLAME [33]. To overcome the lack of large-scale datasets with paired images and 3D models, recent trends have shifted towards self-supervised learning methods [22, 10, 47, 13, 42], using landmark re-projection and/or photometric error. These approaches can yield suboptimal results when landmarks are missing, or color information is compromised due to occlusions. Alternatively, MICA [55] uses a small dataset of paired 2D-3D data to map an image embedding from a face recognition network [11] to the FLAME shape parameters. This technique achieves state-of-the-art results for neutral face shape reconstruction but does not support expressive faces.
@@
A few works have specifically addressed the problem of reconstructing faces under occlusion. Egger et al. [18, 19] proposed a probabilistic optimization approach that simultaneously solves for model parameters and segmentation regions using an expectation–maximization approach. FOCUS [31] follows a similar idea within a self-supervised learning strategy. These methods prioritize obtaining valid reconstructions for the non-occluded regions but do not address the ambiguity of the problem, which requires a distribution of solutions. Recently, Diverse3D [14] tackled this by employing a mesh-based variational autoencoder (VAE) for shape completion, combined with Determinantal Point Process (DPP)  [29] for sampling diverse solutions in expression space. However, the method often results in unrealistic and exaggerated reconstructions (see Fig. 2) because DDP is not designed to capture the statistical properties of the data.
@@
Ranking plays an essential role in Information Retrieval [35] commonly used for document search and retrieval [28, 25], image search [38], and recommendation systems [36]. Some recent works focusing on text ranking with large language models, such as T5 [54], incorporate ranking losses as a fine-tuning tool to optimize model performance. RankGAN [34] uses ranking as a rewarding mechanism to train a language generator to produce higher-quality descriptions. To the best of our knowledge, ranking has not been explored in the context of diffusion models to select an optimal sample.
3 Preliminaries
3.1 FLAME Parametric Model
FLAME [33] is a parametric 3D face model combining separate shape and expression spaces to form a 3D head mesh $M$ with $n=5023$ vertices. Given the shape $\beta\in\mathbb{R}^{|\beta|}$, expression $\psi\in\mathbb{R}^{|\psi|}$, and pose $\theta\in\mathbb{R}^{|\theta|}$ parameters, FLAME outputs a mesh $M$ as $M(\beta,\theta,\psi)=\text{LBS}\big{(}{\cal T}_{p}(\beta,\theta,\psi),{\cal J}(\beta),\theta,{\cal W}\big{)}$ , where LBS is the linear blend skinning function using weights ${\cal W}\in\mathbb{R}^{4\times n}$ and joint regressor ${\cal J}(\mbox{\boldmath$\beta$})$, and $T_{p}(\beta,\theta,\psi)\!={\cal T}+{\cal B}_{S}(\beta,{\cal S})+{\cal B}_{P}(\theta,{\cal P})+{\cal B}_{E}(\psi,{\cal E})$ deforms a template mesh into the desired pose, with $B_{S},B_{E},B_{P}$ the shape, expression and pose-corrective blendshapes, respectively. In the original paper FLAME uses $|\beta|=300$ for shape, and $|\psi|=100$ dimensions for expression, which we follow here.
3.2 Information Retrieval (IR) Ranking
Ranking is an important technique used in the information retrieval domain to retrieve relevant documents given a query. There are three primary learning-based relevance ranking techniques: pointwise, pairwise and listwise ranking. The pointwise model takes in a single $q)$-document($d_{i}$) pair and gives the relevance score($s_{i}$): $\bar{r}(q,d_{i})=s_{i}$, while pairwise ranking orders the two documents based on importance and relevance to the query: $\bar{r}(q,d_{i},d_{j})=P(d_{i}\triangleright d_{j})$. Unlike other methods, listwise ranking [40] takes into account all the documents and ranks them by calculating the importance score of each document: $\bar{r}(q,d_{i},\ldots,d_{n})=(r_{1},\ldots,r_{n})$, which is the method we adopt in our network.
@@ 4 Method
Our method takes as input a single-view image of an occluded face and generates a set of 3D faces as output. The goal is to produce reconstructions that explore a diverse range of expressions in the occluded areas while accurately capturing the visible regions of the input image.
@@
Figure 3: Architecture of OFER. Given an input image, the Identity Generative Network samples N shape parameters and Expression Generative Network samples N expression parameters. The reconstructed shape is passed to Identity Ranking Network to select one of the optimal shapes. The selected shape parameters is combined with expression parameters to output the diverse expressive face reconstructions.
@@
The architecture of OFER consists of three components. First, the Identity Generative Network (Fig. 4), a DDPM conditional diffusion model, outputs a set of FLAME  [33] shape coefficients. This step recovers a distribution of plausible neutral (expressionless) 3D faces. Next, the Identity Ranking Network (Fig. 5), a small MLP, evaluates and ranks the shape samples given by the previous network. This step selects one of the unique shape coefficients that best explains the image. Finally, the Expression Generative Network (Fig. 6), another diffusion-based network, generates a diverse set of FLAME expression coefficients. All three are conditioned on the same input image. By combining the selected shape coefficient with the set of expression hypotheses, we obtain the final set of 3D reconstructions. The overall architecture of OFER is shown in  Fig. 3. We detail each of its components in the following sections.
@@ 4.1 Identity Generative Network (IdGen)
Figure 4: Architecture of the Identity Generative Network. During training, the input image is encoded into a 512-dimensional embedding using the ArcFace [11] and is used as a condition for the 3-layer 1D U-Net Diffusion Network. The ground truth 300-dimensional 3DMM shape coefficients of the image serve as input to the network.
We first train the shape network, IdGen, to produce FLAME shape coefficients $S\in\mathbb{R}^{300}$. This network is a 1D U-Net  [43] with self-attention layers that diffuses $S$ into unit Gaussian noise $\varepsilon\sim\mathcal{N}(0,I_{300})$ through time-steps $t={0..T}$. The input image is used as conditional guidance by encoding it into a 512-dimensional feature vector $c_{a}\in\mathbb{R}^{512}$ using the ArcFace [11] encoder, which is trained with angular margin loss. We adopt the usage of ArcFace from MICA [55] for its highly discriminative features in face recognition. Following MICA [55], we freeze all layers of the encoder except the last, allowing the fine-tuning of the encoded features during training. As commonly done with diffusion models, we employ a sinusoidal embedding, $c_{t}$, to represent the progression of time. During training, we use conditional biasing by concatenating the encoded image conditional feature with each time sequence embedding, $\boldsymbol{C_{\textit{S}}}=(c_{a}||c_{t})$, and passing it to every layer of the network. During sampling, the network initializes from a unit Gaussian noise $\varepsilon\in\mathbb{R}^{300}$, at $t=T$ and iteratively denoises it towards FLAME shape coefficients, $s_{0}=\hat{S}\in\mathbb{R}^{300}$. The network optimizes parameters $\theta_{S}$ by minimizing the following loss function $L_{\varepsilon}$:
$L_{\varepsilon}(\boldsymbol{\theta_{S}})=\mathbb{E}_{t}\Bigl{[}\mathbb{E}_{s_{0},\varepsilon}[|\hat{\varepsilon}_{t}-\varepsilon_{t}|]\Bigr{]}$ (1)
Due to the inherent nature of diffusion models, each generated sample, $\hat{S_{i}}$, exhibits plausible variations in occluded regions, capturing the variance in the input data while preserving the non-occluded facial structure based on the visible cues from the conditioned image.

### v2

@@ Abstract
Reconstructing 3D face models from a single image is an inherently ill-posed problem, which becomes even more challenging in the presence of occlusions. In addition to fewer available observations, occlusions introduce an extra source of ambiguity where multiple reconstructions can be equally valid. Despite the ubiquity of the problem, very few methods address its multi-hypothesis nature. In this paper we introduce OFER, a novel approach for single-image 3D face reconstruction that can generate plausible, diverse, and expressive 3D faces, even under strong occlusions. Specifically, we train two diffusion models to generate the shape and expression coefficients of a face parametric model, conditioned on the input image. This approach captures the multi-modal nature of the problem, generating a distribution of solutions as output. However, to maintain consistency across diverse expressions, the challenge is to select the best matching shape. To achieve this, we propose a novel ranking mechanism that sorts the outputs of the shape diffusion network based on predicted shape accuracy scores. We evaluate our method using standard benchmarks and introduce CO-545, a new protocol and dataset designed to assess the accuracy of expressive faces under occlusion. Our results show improved performance over occlusion-based methods, while also enabling the generation of diverse expressions for a given image.
Figure 1: Reconstructions generated by OFER. Our method can reconstruct faces from a single image under hard occlusions (a), providing multiple solutions with diverse expressions that align with the input image (d). We use two diffusion models that denoise shape and expression parameters of FLAME conditioned on the image. A novel ranking mechanism selects an optimal identity (c) from the generated set of shapes, on top of which the expression variants are applied to obtain the final results.
11footnotetext: Project website: https://ofer.is.tue.mpg.de/
@@ 1 Introduction
The problem becomes even more challenging when images are subjected to occlusions, as shown in Fig. 1. The main difficulty arises from the face being only partially visible, introducing an additional source of ambiguity: the occluded areas can now correspond to an infinite number of valid shapes, making it a multi-hypothesis reconstruction problem. Occlusions are very common in images captured in uncontrolled environments due to factors such as hair, accessories, medical masks, or even strong profile poses and head rotations (e.g. Fig. 9). Despite its prevalence, reconstruction under such conditions has seldom been addressed.
@@
3D face reconstruction typically involves recovering the parameters of a statistical 3D facial model, such as the 3D Morphable Model (3DMM) [4], either via model fitting [4, 1, 24] or regression [9, 12, 22, 51]. While these works may handle milder occlusions depending on their training data, their performance deteriorates significantly under more severe obstructions (see e.g. Fig. 2). A few methods have been specifically proposed to address the challenge of occluded faces [17, 18, 31, 13]. Most offer a unique solution due to their deterministic nature, ignoring the multi-hypothesis aspect of the problem [17, 18, 31]. An exception to this is Diverse3D [13], which employs Determinantal Point Process (DPP) [29] to sample a diverse set of outputs. While DDP is designed to capture the diversity of the data, it fails to adequately represent the underlying distribution of facial geometry, often resulting in unrealistic and exaggerated reconstructions as shown in Fig. 2.
@@
To ensure plausible face reconstructions while addressing the challenges of diverse sampling, we introduce OFER, a novel method for reconstructing 3D faces under occlusions. At its core, OFER employs two denoising diffusion probabilistic models (DDPMs) [26] to generate the shape and expression coefficients of the FLAME [32] parametric face model, conditioned on the input image. The ability of diffusion models to learn the underlying data distribution enables OFER to produce multiple plausible hypotheses. We leverage two pre-trained face image recognition networks as image encoders, and following MICA [59], train our models using only a small dataset of paired 2D-3D data.
Generating diverse solutions is crucial for reconstructing faces under occlusions. Equally important is selecting a unique, consistent shape that best represents the face in the image across all generated expressions. This is achievable since identity is generally better defined under occlusions.
@@
To achieve this, we propose a novel ranking mechanism that evaluates and ranks the samples generated by the shape diffusion network, effectively filtering out poor quality samples. Specifically, given an input image along with N generated shape samples, the ranking network scores and ranks the candidates, selecting the top-ranked one. To the best of our knowledge, such a ranking mechanism as selection process for diffusion models has not been considered before.
@@
Figure 2: Comparison against expression reconstruction methods. We show results from EMOCA [9] (pink); samples generated by Diverse3D [13] (blue); and samples generated by our method (green). EMOCA can only reconstruct a single solution due to its deterministic nature, while Diverse3D shows non-plausible faces. OFER (our method) generates diverse 3D faces with plausible expressions.
@@
To evaluate our approach, we introduce a new dataset and evaluation protocol, CO-545, derived from the CoMA dataset [42]. CO-545 includes 545 ground-truth pairs of occluded images and corresponding FLAME non-occluded 3D vertices. Experimental results show that our method effectively generates multiple hypotheses for single in-the-wild occluded images, achieving superior quality and diversity on CO-545. Additionally, when evaluated on the NoW benchmark [47], our method demonstrates improved performance with the ranking-selection mechanism.
@@ 2 Related Work
3D face reconstruction from a single image has been a key research area for several decades [19], with approaches broadly categorized into model-free [11, 21, 45, 16, 2, 27, 48, 55] and model-based methods [59, 22, 9, 56, 43]. Model-free approaches estimate 3D geometry directly from images, while model-based ones recover the low-dimensional parameters of a statistical model of the 3D face such as BFM [39] or FLAME [32]. To overcome the lack of large-scale datasets with paired images and 3D models, recent trends have shifted towards self-supervised learning [22, 9, 50, 12, 43], using landmark re-projection and/or photometric error. These works can yield suboptimal results when landmarks are missing or when color information is compromised due to occlusions. Alternatively, MICA [59] uses a small dataset of paired 2D-3D data to map an image embedding from a face recognition network [10] to the FLAME shape parameters. This technique achieves state-of-the-art results for neutral face shape reconstruction but does not support expressive faces.
@@
A few works have specifically addressed the problem of reconstructing faces under occlusion. Egger et al. [17, 18] proposed a probabilistic optimization approach that simultaneously solves for model parameters and segmentation regions using an expectation–maximization approach. FOCUS [31] follows a similar idea within a self-supervised learning strategy. These methods prioritize obtaining valid reconstructions for the non-occluded regions but do not address the ambiguity of the problem, which requires a distribution of solutions. Recently, Diverse3D [13] tackled this by employing a mesh-based variational autoencoder (VAE) for shape completion, combined with Determinantal Point Process (DPP)  [29] for sampling diverse solutions in expression space. However, the method often results in unrealistic and exaggerated reconstructions (see Fig. 2), since DDP is not designed to capture the statistical properties of the data.
@@
Ranking plays an essential role in Information Retrieval [34], commonly used for sorting documents by relevance [28, 25], image search [38], and recommendation systems [35]. Learning to Rank methods [34] perform this task by making use of supervised machine learning. RankGAN [33] uses ranking as a rewarding mechanism to train a language generator to produce higher-quality descriptions. Some recent works such as T5 [58] focus on text ranking with large language models, and incorporate ranking losses as a fine-tuning tool to optimize model performance. To the best of our knowledge, ranking has not been explored in the context of diffusion models to select an optimal sample.
@@ 3 Method
Our method takes as input a single-view image of an occluded face and generates a set of 3D faces as output. The goal is to produce reconstructions that explore a diverse range of expressions in the occluded areas, while accurately capturing the visible regions of the input image.
@@
Figure 3: Overview of OFER. Given an input image, the Identity Generative Network (IdGen) samples N shape parameters. The reconstructed shapes are then passed to the Identity Ranking Network (IdRank) to select a unique identity. Finally, the Expression Generative Network (ExpGen) generates N expression parameters, which are combined with the selected shape to output diverse and expressive face reconstructions (bottom row).
@@
OFER is structured around three key components. First, the Identity Generative Network (IdGen, Fig. 4 and Sec. 3.2), a DDPM conditional diffusion model, outputs a set of FLAME  [32] shape coefficients, recovering a distribution of plausible neutral 3D faces. Next, the Identity Ranking Network (IdRank, Fig. 5 and Sec. 3.3), a small MLP, evaluates and ranks the shape samples given by the previous network. This step selects a unique shape coefficient that best explains the image. Finally, the Expression Generative Network (ExpGen, Fig. 4 and Sec. 3.4), another conditional DDPM network, generates a diverse set of FLAME expression coefficients. We use default DDPM framework [26] with DDPM sampler with 1000 sampling time steps for both IdGen and ExpGen. The three networks are conditioned on the same input image. By combining the selected shape coefficient with the set of expression hypotheses, we obtain the final set of 3D reconstructions. The overall architecture of OFER is shown in  Fig. 3. We detail each of its components in the following sections.
3.1 Preliminaries
FLAME Model.
FLAME [32] is a parametric 3D face model combining separate shape and expression spaces to form a 3D head mesh $M$ with $n=5023$ vertices. Given the shape $S\in\mathbb{R}^{|S|}$, expression $E\in\mathbb{R}^{|E|}$, and pose $P\in\mathbb{R}^{|P|}$ parameters, FLAME produces a mesh $M$ as
$M(S,E,P)=\text{LBS}\big{(}{\cal T}(S,E,P),{\cal J}(S),P,{\cal W}\big{)},$ (1)
where LBS is the linear blend skinning function using weights ${\cal W}\in\mathbb{R}^{4\times n}$ and joint regressor ${\cal J}(S)\in\mathbb{R}^{3K}$, and $T(S,E,P)\!={\cal T}+{\cal B}_{S}(S)+{\cal B}_{E}(E)+{\cal B}_{P}(P)$ deforms a template mesh ${\cal T}$ using the shape ${\cal B}_{S}$, expression ${\cal B}_{E}$ and pose-corrective ${\cal B}_{P}$ blendshapes. FLAME uses $|S|=300$ dimensions for shape and $|E|=100$ dimensions for expression.
Denoising Diffusion Probabilistic Model (DDPM).
The underlying mechanism of DDPM is the transformation of an unknown data distribution, $p(X)\in\mathcal{R}^{d}$ into a simple known distribution. Here, $X$ represents the data from the underlying distribution. This transformation is achieved by iteratively applying a transition kernel $q$ via a Markov chain process with infinitesimal time steps, ensuring a stationary distribution at each time step $t$ i.e $x_{t}\sim q(x_{t}|x_{t-1}),\forall t>0$ where $x\in X$. DDPM models these transformations by parameterizing a neural network to captures complex dependencies in the data by modeling the sequential evolution of the data distribution. In the work of Sohl et al.[49], the known distribution is set as gaussian with a decaying variance schedule $\beta_{t}\in\mathcal{R}$ such that $q(x_{t}|x_{t-1}=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}x_{t-1},\beta_{t}I);q(x_{T})=\mathcal{N}(x_{T};0,I)$ due to its simplicity and tractability. Denoising in DDPM involves reversing this process by starting with unit gaussian noise and attempting to model the underlying data distribution $p^{\prime}(X)$.
Learning-based Ranking.
Ranking is an important technique used in the information retrieval domain to retrieve relevant documents given a query. There are three primary learning-based ranking techniques [34]: pointwise, pairwise and listwise. The pointwise model takes in a single $q)$-document($d_{i}$) pair and gives the relevance score ($s_{i}$) of each in isolation, which requires the ground-truth score: $\bar{r}(q,d_{i})=s_{i}$. Pairwise ranking methods compare two documents against each other based on importance and relevance to the query: $\bar{r}(q,d_{i},d_{j})=P(d_{i}\triangleright d_{j})$. Finally, listwise ranking [41] generates an optimal order of a list of documents by calculating the importance score of each: $\bar{r}(q,d_{i},\ldots,d_{n})=(r_{1},\ldots,r_{n})$, which is the method we adopt in our network.
@@ 3.2 Identity Generative Network (IdGen)
Figure 4: Overview of the identity and expression generative networks (IdGen, in blue, and ExpGen, in green). For IdGen, the input image is encoded into a 512-dimensional embedding using ArcFace [10]. This serves as a condition for the 1D U-Net diffusion network, which is trained to denoise 300-dimensional noise into FLAME shape coefficients, S. For ExpGen, the input image is encoded into a 1024-dimensional embedding using the FaRL [57] and ArcFace [10] encoders. The embedding serves as a condition for the 1D U-Net diffusion network, which is trained to denoise 50-dimensional noise into FLAME expression coefficients, E.
[…]
The input image is used as condition by encoding it into a 512-dimensional feature vector $c_{a}\in\mathbb{R}^{512}$ using the ArcFace [10] encoder. Following MICA [59], we freeze all layers of the ArcFace encoder except the last, such that the encoded features can be fine-tuned during training. During sampling, we generate a set of $N$ noise vectors $\{s_{T}^{i}\in\mathbb{R}^{300}\}_{i=1}^{N}$ and iteratively denoise them using $\epsilon^{S}_{\theta}$ to obtain the FLAME shape coefficients, $\{s_{0}^{i}=\hat{S}_{i}\in\mathbb{R}^{300}\}_{i=1}^{N}$. We optimize the parameters $\theta$ by minimizing the following loss function $L_{\varepsilon}$:
$L_{\varepsilon}(\theta)=\mathbb{E}_{t}\Bigl{[}\mathbb{E}_{s_{0},\varepsilon}[|\epsilon^{S}_{\theta}(s_{t})-\varepsilon_{t}|]\Bigr{]}.$ (2)
where $\varepsilon_{t}$ is fixed Gaussian noise with variance proportional to the linear time step $t$ [26]. Due to the nature of diffusion models, each generated sample $\hat{S_{i}}$ exhibits plausible variations in the occluded regions, capturing the variance in the input data.